Determine projau the orthogonal projection of u and a and deduce ||projau|| for

a.) u=<-1 3>, a=<-1,-3>;

b.)u=<-2,1,-3>, a=<-2,1,2>.


1
Expert's answer
2021-06-24T16:06:32-0400
"proj_{\vec a}\vec u=\dfrac{\vec u\cdot\vec a}{|\vec a|^2}\vec a"

(2.1)


"\vec u=\langle-1,3\rangle, \vec a=\langle-1,-3\rangle"

"\vec u\cdot\vec a=-1(-1)+3(-3)=-8"

"|\vec a|^2=(-1)^2+(-3)^2=10"

"proj_{\vec a}\vec u=\dfrac{\vec u\cdot\vec a}{|\vec a|^2}\vec a=\dfrac{-8}{10}\langle-1,-3\rangle"

"=\langle0.8,-0.3\rangle""|proj_{\vec a}\vec u|=\sqrt{(0.8)^2+(-0.3)^2}=\sqrt{0.73}"


(2.2)


"\vec u=\langle-2,1, -3\rangle, \vec a=\langle-2,1, 2\rangle"

"\vec u\cdot\vec a=-2(-2)+1(1)-3(2)=-1"

"|\vec a|^2=(-2)^2+(1)^2+(2)^2=9"

"proj_{\vec a}\vec u=\dfrac{\vec u\cdot\vec a}{|\vec a|^2}\vec a=\dfrac{-1}{9}\langle-2,1,2\rangle"

"=\langle\dfrac{2}{9},-\dfrac{1}{9},-\dfrac{2}{9}\rangle""|proj_{\vec a}\vec u|=\sqrt{(\dfrac{2}{9})^2+(-\dfrac{1}{9})^2+(-\dfrac{2}{9})^2}=\dfrac{1}{3}"